# Rectified Point Flow: Generic Point Cloud Pose Estimation

### v1

@@ Abstract
We present Rectified Point Flow, a unified parameterization that formulates pairwise point cloud registration and multi-part shape assembly as a single conditional generative problem. Given unposed point clouds, our method learns a continuous point-wise velocity field that transports noisy points toward their target positions, from which part poses are recovered. In contrast to prior work that regresses part-wise poses with ad-hoc symmetry handling, our method intrinsically learns assembly symmetries without symmetry labels. Together with a self-supervised encoder focused on overlapping points, Rectified Point Flow achieves a new state-of-the-art performance on six benchmarks spanning pairwise registration and shape assembly. Notably, our unified formulation enables effective joint training on diverse datasets, facilitating the learning of shared geometric priors and consequently boosting accuracy. Our code and models are available at https://rectified-pointflow.github.io/.
@@ 1 Introduction
In this work, we revisit 3D pose regression and propose a conditional generative model for generic point cloud pose estimation that casts the problem as learning a continuous point-wise flow over the input geometry, effectively capturing priors over assembled shapes. More specifically, our method, Rectified Point Flow, models the motion of points from random Gaussian noise in Euclidean space toward the point clouds of assembled objects, conditioned on the unposed part point clouds. This learned flow implicitly encodes part-level transformations, enabling both discriminative pose estimation and generative shape assembly within a single framework.
To further instill geometric awareness of inter-part relationships, we pretrain the encoder of the conditional point clouds on large-scale 3D shape datasets using a self-supervised task: predicting point-wise overlap across parts, formulated as a binary classification task. While GARF [11] also highlights the value of encoder pretraining for a flow model, it relies on mesh-based physical simulation [12] to generate fracture-based supervision signals. In contrast, we introduce a lightweight and scalable alternative that constructs pretraining data by computing geometric overlap between parts. Our data generation is agnostic to data sources tailored for different tasks—including part segmentation [13, 14, 15], shape assembly [12, 16, 17], and registration [18, 19]—without requiring watertight mesh or simulation, an important step towards scalable pretraining for pose estimation.
Our flow-based pose estimation departs from traditional pose-vector regression in three ways: (i) Geometric grounding: Rather than regressing poses directly in $\mathrm{SE}(3)$, we operate in Euclidean space over dense point clouds. This makes the model inherently robust to symmetries, part interchangeability, and spatial ambiguities that often challenge conventional methods; (ii) Joint shape-pose reasoning: By training to predict the final assembled point cloud, our model learns to reconstruct the shape of the object; and (iii) Scalable shape prior: We cast registration and assembly task as “reconstruct the completed assembly.” This unified objective lets a single network learn from heterogeneous datasets, yielding scalable training and transferable geometric knowledge across standard pairwise registration, fracture reassembly, and complex furniture assembly tasks.
Our main contributions are summarized as follows:
  • We propose Rectified Point Flow, a conditional generative model for generic point cloud pose estimation that addresses both pairwise registration and multi-part assembly tasks and achieves state-of-the-art performances on all the tasks.
  • We propose a generalizable pretraining strategy for learning geometric awareness over inter-part relationships across several 3D shape datasets, and formulate it as point-wise overlap prediction across parts.
@@ 2 Related Work
Euler angles and quaternions are the predominant parametrization of rotation in various pose regression tasks [20, 21, 22, 23, 24, 25, 11, 26] due to their simplicity and usability. As Euler angles and quaternions are discontinuous representations, Zhou et al. [27] proposed to represent 3D rotation with a continuous representation for neural networks using 5D and 6D vectors. In contrast to directly regressing pose vectors, other methods train networks to find sparse correspondences between image pairs or point cloud pairs and extract pose vectors using Singular Value Decomposition (SVD) [28, 29, 8, 30, 31, 32]. More recently, RayDiffusion [33] proposed to represent camera poses as ray bundles, naturally suited for coupling image features and transformer architectures. DUSt3R [34] directly regresses the pointmap of each camera in a global reference frame and then extracts the camera pose using RANSAC-PnP [35, 36]. Our proposed rectified point flow, extends the pointmap representation for learning generalizable pose estimation on point cloud registration and assembly.
@@
Multi-part registration and shape assembly generalize pairwise relative pose estimation to multiple parts, with applications in furniture assembly [17] and shape reassembly [12]. Methods [43, 44, 26, 45, 46] tackle the multi-part registration problem by estimating the transformation for each rigid part in the scene (multi-source and multi-target). Multi-part shape assembly differs as a task from registration because it has multi-source input and a canonical target, and each part has almost ‘zero’ overlap with each other. Chen et al. [47] adopt an adversarial learning scheme to examine the plausibility for different shape configurations. Wu et al. [48] leverage $\mathrm{SE(3)}$-equivariant representation to handle pose variations in shape assembly. DiffAssembly [49] and PuzzleFusion [24, 25] leverage diffusion models to predict the transformation for each part. GARF [11] combines fracture-aware pretraining with a flow matching model to predict per-part transformation. However, these approaches do not explicitly address part symmetry or interchangeability as effectively as our method. A concurrent work [50] tackles multi-part assembly via equivariant flow matching, relying on an $\mathrm{SE}(3)$-equivariant network and carefully designed flow trajectories to account for part symmetry; in contrast, we handle the part symmetry with a simpler straight-line flow formulation in Euclidean space. Moreover, Rectified Point Flow is the first class-agnostic solution for furniture assembly on the PartNet [15] and IKEA-Manual [17] datasets.
Figure 1: Pose Estimation via Rectified Point Flow. Our formulation supports shape assembly and pairwise registration tasks in a single framework. Given a set of unposed part point clouds $\{\bar{\bm{X}}_{i}\}_{i\in\Omega}$, Rectified Point Flow predicts each part’s point cloud at the target assembled state $\{\hat{\bm{X}}_{i}{(0)}\}_{i\in\Omega}$. Subsequently, we solve Procrustes problem via SVD between the condition point cloud $\bar{\bm{X}}_{i}$ and the estimated point cloud $\hat{\bm{X}}_{i}(0)$ to recover the rigid transformation $\hat{T}_{i}$ for each non-anchored part.
@@ 3 Pose Estimation via Rectified Point Flow
Rectified Point Flow addresses the multi-part point cloud pose estimation problem, defined in Sec. 3.1. The overall pipeline consists of two consecutive stages: self-supervised overlap-aware point encoding (Sec. 3.2) and conditional Rectified Point Flow (Sec. 3.3). Finally, we explain how our formulation inherently addresses the challenges posed by symmetric and interchangeable parts in Sec. E.
@@ 3.2 Overlap-aware Point Encoding
Pose estimation relies on geometric cues from mutually overlapping regions among connected parts [29, 32, 11]. In our work, we address this challenge through a self-supervised pretraining approach that develops a task-agnostic, overlap-aware encoder capable of producing pose-invariant point features. As illustrated in Fig. 2, we train an encoder $F$ to identify overlapping points in different parts. Given a set of unposed parts $\{\bm{X}_{i}\}_{i\in\Omega}$, we first apply random rigid transforms $\tilde{T}_{i}\in\mathrm{SE}(3)$ and compose transformed point clouds $\tilde{\bm{X}}_{i}=\tilde{T}_{i}\bm{X}_{i}$ as input to the encoder. These data augmentations enable the encoder to learn more robust pose-invariant features. The encoder then computes per-point features $C_{i,j}\in\mathbb{R}^{d}$ for the $j$-th point on part $i$, after which an MLP overlap prediction head estimates the overlap probability $\hat{p}_{i,j}$. The binary ground-truth label $p_{i,j}$ is 1 if point $\tilde{\mathbf{x}}_{i,j}$ falls within radius $\epsilon$ of at least one point in other parts.
We train both the encoder and the overlap head using binary cross-entropy loss. For objects without predefined part segmentation, we employ off-the-shelf 3D part segmentation methods to generate the necessary labels. The features extracted by our trained encoder subsequently serve as conditioning input for our Rectified Point Flow model.
@@ 3.3 Generative Modeling for Pose Estimation
The overlap-aware encoder identifies potential overlap regions between parts but cannot determine their final alignment, particularly in symmetric objects that allow multiple valid assembly configurations. To address this limitation, we formulate the point cloud pose estimation as a conditional generation task. With this approach, Rectified Point Flow leverages the extracted point features to sample from the conditional distribution of all feasible assembled states across multi-part point clouds, generating estimates that maximize the likelihood of the conditional input point cloud. By recasting pose estimation as a generative problem, we naturally accommodate the inherent ambiguities arising from symmetry and part interchangeability in the data.
@@ 3.4 Invariance Under Rotational Symmetry and Interchangeability
Theorem 1 ($\mathcal{G}$-invariance of the learning objective).
@@ 4 Experiments
We use PointTransformerV3 (PTv3) [54] as the backbone for point cloud encoder, and use Diffusion Transformer (DiT) [55] as our flow model. Each DiT layer applies two self-attention stages: (i) part-wise attention to consolidate part-awareness, and (ii) global attention over all part tokens to fuse information. We stabilize the attention computation by applying RMS Normalization [56, 57] to the query and key vectors per head before attention operations. We sample the time steps from a U-shaped distribution following [58]. We pre-train the PTv3 encoder on all datasets with an additional subset of Objaverse [14] meshes, where we apply PartField [13] to obtain annotations. After pretraining, we freeze the weights of the encoder. We train our flow model on 8 NVIDIA A100 80GB GPUs for 400k iterations with an effective batch size of 256. We use the AdamW [59] optimizer with an initial learning rate $5\times 10^{-4}$ which is halved every 25k iterations after the first 275k iterations.
Table 1: Dataset Statistics. We train our model on seven datasets with varying sizes and complexity.
Dataset Task Train & Val Test
@@
IKEA-Manual [17] Assembly 84 [2, 19] 18 [2, 19]
TwoByTwo [16] Assembly 308 [2, 2] 144 [2, 2]
PartNet-Assembly Assembly 23755 [2, 64] 261 [2, 64]
BreakingBad [12] Assembly 35114 [2, 49] 265 [2, 49]
TUD-L [18] Registration 19138 [2, 2] 300 [2, 2]
ModelNet 40 [19] Registration 19680 [2, 2] 260 [2, 2]
Objaverse [14] Pre-training Only 63199 [3, 12] 6794 [3, 12]
@@ 4.1 Experimental Setting
For the multi-part shape assembly task, we experiment on the BreakingBad [12], TwoByTwo [16], PartNet [15], and IKEA-Manual [17] datasets. The PartNet dataset has been processed for the shape assembly task following the same procedure as [17] but includes all object categories; we refer to this version as PartNet-Assembly. Evaluation of the pairwise registration is performed on the TUDL [18] and ModelNet 40 [19] datasets. We split all datasets into train/val/test sets following existing literature for fair comparisons. The statistics of all datasets are in Tab. 1.
@@
We evaluate our method against state-of-the-art methods for pairwise registration and shape assembly. For pairwise registration, we compare against DCPNet [31], RPMNet [30], GeoTransformer [32], and Diff-RPMNet [22]. For shape assembly, we compare against MSN [47], SE(3)-Assembly [48], Jigsaw [60], PuzzleFussion++ [25], and GARF [11]. We did not include the results of [50] because it is evaluated only on samples up to 8 parts (in BreakingBad-Everyday) and their models and codes have not yet been released. We report our performance in two training configurations: dataset-specific training in which models are trained independently for each dataset (Ours (Single)), and joint training in which a single model is trained on all datasets (Ours (Joint)).
@@ 4.2 Evaluation
We report pose accuracy for shape assembly and pairwise registration in Tab. 2  [^1] We found that the BreakingBad benchmark [12, 11, 25] originally computed rotation error (RE) using the RMSE of Euler angles, which is not a proper metric on $\mathrm{SO(3)}$. To ensure consistency, we re-evaluate GARF using the geodesic distance between rotation matrices via the Rodrigues formula [61, 31, 29, 32]. and Tab. 3, respectively. Our model outperforms all existing approaches by a substantial margin. For multi-part assembly, the closest competitor is GARF [11], which formulates per-part pose estimation as 6-DoF pose regression; see Figs. 4 and 5. We attribute our superior results to two key advantages of Rectified Point Flow: (i) in contrast to our closest competitor GARF [11] which performs 6-DoF pose regression, our dense shape-and-pose parametrization helps the model learn better global shape prior and fine-grained geometric details more effectively; and (ii) our generative formulation natively handles part symmetries and interchangeability. For pairwise registration, GARF–despite being retrained on target datasets–fails to generalize beyond the original task. In contrast, our method achieves a new state-of-the-art performance on registration benchmarks, outperforming methods designed solely for registration (e.g., GeoTransformer [32] and Diff-RPMNet [22]) and demonstrating strong generalization across different datasets (Fig 5). We also achieve the strongest shape assembly performance on IKEA-Manual [17]; for more details on evaluation and visualizations, see supplementary.
@@
Joint Training.
By recasting pairwise registration as a two-part assembly task, our unified formulation enables joint training on all six datasets—including very small sets like TwoByTwo (308 samples) and IKEA-Manual (84 samples). Ours (Joint) consistently matches or outperforms the dataset-specific (Ours (Single)) models. For example, on TwoByTwo the rotation error drops from 18.7 to 13.2 ($\approx$30%), and on BreakingBad from 9.6 to 7.4 ($\approx$23%), while on ModelNet40, the rotation error is reduced from 1.37° to 0.93°. These results demonstrate that joint training enables the model to learn shared geometric priors—such as part symmetries, common pose distributions, and cross-dataset correlations—which substantially boosts performance particularly on datasets with limited training samples.
@@
We demonstrate our model’s ability to handle symmetry (Sec. E) on IKEA-Manual [17], a dataset with symmetric assembly configurations. As shown in Fig. 7, while being only trained on a single configuration (left), Rectified Point Flow samples various reasonable assembly configurations (right), conditioned on the same input unposed point clouds. Note how our model permutes the 12 repetitive vertical columns and swaps the two middle baskets, yet always retains the non-interchangeable top and footed bottom baskets in their unique positions.
@@
BreakingBad-Everyday [12]
Encoder Pre-training Task RE ↓ TE ↓ Part Acc ↑
@@
MLP No Pre-training 41.7 12.3 68.3
PTv3 No Pre-training 18.5 4.9 79.5
PTv3 Instance Segmentation 16.7 4.4 80.9
PTv3 Overlap Detection (ours) 9.6 1.8 93.5
Ablation on Self-supervised Pretraining.
Tab. 4 compares four pretraining strategies for our flow-based assembly model on BreakingBad–Everyday [12]. The first two encoders (MLP and PTv3 without pre-training) are trained jointly with the flow model. The last two encoders are PTv3 pretrained on instance segmentation and our overlap-aware prediction tasks, respectively. Their pretrained weights are frozen during flow model training. We find that PTv3 is a more powerful feature encoder compared to the MLP, and pretraining on instance segmentation can already extract useful features for pose estimation, while our proposed overlap-aware pretraining leads to the best accuracy. We hypothesize that, although the segmentation backbone provides strong semantic features, only our overlap prediction task explicitly encourages the encoder to learn fine-grained part interactions and pre-assembly cues—critical for precise assembly and registration.

### v2

@@ Abstract
We present Rectified Point Flow, a unified parameterization that formulates pairwise point cloud registration and multi-part shape assembly as a single conditional generative problem. Given unposed point clouds, our method learns a continuous point-wise velocity field that transports noisy points toward their target positions, from which part poses are recovered. In contrast to prior work that regresses part-wise poses with ad-hoc symmetry handling, our method intrinsically learns assembly symmetries without symmetry labels. Together with an overlap-aware encoder focused on inter-part contacts, Rectified Point Flow achieves a new state-of-the-art performance on six benchmarks spanning pairwise registration and shape assembly. Notably, our unified formulation enables effective joint training on diverse datasets, facilitating the learning of shared geometric priors and consequently boosting accuracy. Our code and models are available at https://rectified-pointflow.github.io/.
@@ 1 Introduction
In this work, we revisit 3D pose regression and propose a generative approach for generic point cloud pose estimation that casts the problem as learning a continuous point-wise flow field over the input geometry, effectively capturing priors over assembled shapes. Our method, Rectified Point Flow, models the motion of points from random Gaussian noise in Euclidean space toward the point clouds of assembled objects. This learned flow implicitly encodes part-level transformations, enabling both discriminative pose estimation and generative shape assembly within a single framework. Rectified Point Flow consists of an encoder that extracts point-wise features, and a flow model that, given the features, predicts the final assembled positions.
To instill geometric awareness of inter-part relationships, we pretrain the encoder on large-scale 3D shape datasets: predicting point-wise overlap across parts, formulated as a binary classification task. While GARF [11] also highlights the value of encoder pretraining for a flow model, it relies on mesh-based physical simulation [12] to generate fracture-based supervision signals. In contrast, we introduce a lightweight and scalable alternative that constructs pretraining data by computing geometric overlap between parts. Our data generation is agnostic to data sources tailored for different tasks—including part segmentation [13, 14, 15], shape assembly [12, 16, 17], and registration [18, 19]—without requiring watertight mesh or simulation, an important step towards scalable pretraining for pose estimation.
Our flow-based pose estimation departs from traditional pose-vector regression in three ways: (i) Joint shape-pose reasoning: We cast the registration and assembly tasks as one unified task that reconstructs the complete shape while simultaneously enabling the estimation of part poses; (ii) Scalable shape prior learning: By training to predict the final assembled point cloud, our model learns from heterogeneous datasets and part definitions, yielding scalable training and transferable geometric knowledge across standard pairwise registration, fracture reassembly, and complex furniture assembly tasks; and (iii) Intrinsic symmetry handling: Rather than regressing pose vectors directly in $\mathrm{SE}(3)$ space, we operate in Euclidean space over dense point clouds. This makes the model inherently robust to symmetries, part interchangeability, and spatial ambiguities that often challenge conventional methods. Our main contributions are summarized as follows:
  • We propose Rectified Point Flow, a generative approach for generic point cloud pose estimation that addresses both pairwise registration and multi-part assembly tasks and achieves state-of-the-art performances on all the tasks.
  • We propose a generalizable pretraining strategy with geometric awareness of inter-part relationships across several 3D shape datasets, and formulate it as point-wise overlap prediction.
@@ 2 Related Work
Euler angles and quaternions are the predominant parametrization of rotation in various pose regression tasks [20, 21, 22, 23, 24, 25, 11, 26] due to their simplicity and usability. As Euler angles and quaternions are discontinuous representations, Zhou et al[27] proposed to represent 3D rotation with a continuous representation for neural networks using 6D and 7D vectors. In contrast to directly regressing pose vectors, other methods train networks to find sparse correspondences between image pairs or point cloud pairs and extract pose vectors using Singular Value Decomposition (SVD) [28, 29, 8, 30, 31, 32]. More recently, RayDiffusion [33] proposed to represent camera poses as ray bundles, naturally suited for coupling image features and transformer architectures. Huang et al[34] adopted a point cloud generative model for policy learning in robot pick-and-place tasks, then recovered relative poses between the object and gripper via SVD. DUSt3R [35] directly regresses the pointmap of each camera in a global reference frame and then extracts the camera pose using RANSAC-PnP [36, 37]. Our proposed Rectified Point Flow, extends the dense point cloud or point map representation for learning generalizable pose estimation on point cloud registration and shape assembly tasks.
@@
Multi-part registration and shape assembly generalize pairwise relative pose estimation to multiple parts, with applications in furniture assembly [17] and shape reassembly [12]. Methods [44, 45, 26, 46, 47] tackle the multi-part registration problem by estimating the transformation for each rigid part in the scene (multi-source and multi-target). Multi-part shape assembly differs as a task from registration because it has multi-source input and a canonical target, and each part has almost ‘zero’ overlap w.r.t. each other. Chen et al. [48] adopt an adversarial learning scheme to examine the plausibility for different shape configurations. Wu et al. [49] leverage $\mathrm{SE(3)}$ equivariant representation to handle pose variations in shape assembly. DiffAssembly [50] and PuzzleFusion [24, 25] leverage diffusion models to predict the transformation for each part. GARF [11] combines fracture-aware pretraining with a flow matching model to predict per-part transformation. These methods, however, do not handle interchangeability and symmetry as well as ours does. Moreover, Rectified Point Flow is the first solution for furniture assembly of 3D shapes on the PartNet-Assembly [15] and IKEA-Manual [17] datasets.
Figure 1: Rectified Point Flow’s Pose-from-Shape Pipeline. Our formulation supports shape assembly (first row) and pairwise registration (second row) tasks in a single framework. Given a set of unposed part point clouds $\{\bar{\bm{X}}_{i}\}_{i\in\Omega}$, Rectified Point Flow predicts each part’s point cloud at the target assembled state $\{\hat{\bm{X}}_{i}{(0)}\}_{i\in\Omega}$. Subsequently, we solve Procrustes problem via SVD between the condition point cloud $\bar{\bm{X}}_{i}$ and the estimated point cloud $\hat{\bm{X}}_{i}(0)$ to recover the rigid transformation $\hat{T}_{i}$ for each non-anchored part.
@@ 3 Pose Estimation via Rectified Point Flow
Rectified Point Flow addresses the multi-part point cloud pose estimation problem, defined in Sec. 3.1. The overall pipeline consists of two consecutive stages: overlap-aware point encoding (Sec. 3.2) and conditional Rectified Point Flow (Sec. 3.3). Finally, we explain how our formulation inherently addresses the challenges posed by symmetric and interchangeable parts in Sec. E.
@@ 3.2 Overlap-aware Point Encoding
Pose estimation relies on geometric cues from mutually overlapping regions among connected parts [29, 32, 11]. In our work, we address this challenge through a pretraining module that develops a task-agnostic, overlap-aware encoder capable of producing pose-invariant point features. As illustrated in Fig. 2, we train an encoder $F$ to identify overlapping points in different parts. Given a set of unposed parts $\{\bm{X}_{i}\}_{i\in\Omega}$, we first apply random rigid transforms $\tilde{T}_{i}\in\mathrm{SE}(3)$ and compose transformed point clouds $\tilde{\bm{X}}_{i}=\tilde{T}_{i}\bm{X}_{i}$ as input to the encoder. These data augmentations enable the encoder to learn more robust pose-invariant features. The encoder then computes per‐point features $C_{i,j}\in\mathbb{R}^{d}$ for the $j$-th point on part $i$, after which an MLP overlap prediction head estimates the overlap probability $\hat{p}_{i,j}$. The binary ground‐truth label $p_{i,j}$ is 1 if point $\tilde{\mathbf{x}}_{i,j}$ falls within radius $\epsilon$ of at least one point in other parts.
We train both the encoder and the overlap head using binary cross‐entropy loss. For objects without predefined part segmentation, we employ off-the-shelf 3D part segmentation methods to generate the necessary labels. The features extracted by our trained encoder subsequently serve as conditioning input for our Rectified Point Flow model.
@@ 3.3 Generative Modeling for Pose Estimation
The overlap‐aware encoder identifies potential overlap regions between parts but cannot determine their final alignment, particularly in symmetric objects that allow multiple valid assembly configurations. To address this limitation, we formulate the point cloud pose estimation as a conditional generation task. With this approach, Rectified Point Flow leverages the extracted point features to sample from the conditional distribution of all feasible assembled states across multi-part point clouds, generating estimates that maximize the likelihood of the conditional input point cloud. By recasting pose estimation as a generative problem, we naturally accommodate the inherent ambiguities arising from symmetry and part interchangeability in the data.
@@ 3.4 Invariance Under Rotational Symmetry and Interchangeability
Theorem 1 ($\mathcal{G}$‑invariance of the learning objective).
@@ 4 Experiments
We use PointTransformerV3 (PTv3) [54] as the backbone for point cloud encoder, and use Diffusion Transformer (DiT) [55] as our flow model. Each DiT layer applies two self‐attention stages: (i) part‐wise attention to consolidate part-awareness, and (ii) global attention over all part tokens to fuse information. We stabilize the attention computation by applying RMS Normalization [56, 57] to the query and key vectors per head before attention operations. We sample the time steps from a U-shaped distribution following [58]. We pre-train the PTv3 encoder on all datasets with an additional subset of Objaverse [14] meshes, where we apply PartField [13] to obtain annotations. After pretraining, we freeze the weights of the encoder. We train our flow model on 8 NVIDIA A100 80GB GPUs for 400k iterations with an effective batch size of 256. We use the AdamW [59] optimizer with an initial learning rate $5\times 10^{-4}$ which is halved every 25k iterations after the first 275k iterations.
Table 1: Dataset statistics. We train our flow model on six datasets with varying sizes, part definitions, and complexities. The encoder is pre-trained on these datasets with an extra Objaverse dataset.
Dataset Task Part Definition Train & Val Test
@@
IKEA-Manual [17] Assembly Reusability and packing 84 [2, 19] 18 [2, 19]
TwoByTwo [16] Assembly Insertable parts 308 [2, 2] 144 [2, 2]
PartNet-Assembly Assembly Semantics and functions 23755 [2, 64] 261 [2, 64]
BreakingBad [12] Assembly Fracture simulation 35114 [2, 49] 265 [2, 49]
TUD-L [18] Registration RGB-D sensor scans 19138 [2, 2] 300 [2, 2]
ModelNet-40 [19] Registration Random partition 19680 [2, 2] 260 [2, 2]
Objaverse 1.0 [14] Pre-training From PartField [13] 63199 [3, 12] 6794 [3, 12]
@@ 4.1 Experimental Setting
For the multi-part shape assembly task, we experiment on the BreakingBad [12], TwoByTwo [16], PartNet [15], and IKEA-Manual [17] datasets. The PartNet dataset has been processed for the shape assembly task following the same procedure as [17] but includes all object categories; we refer to this version as PartNet-Assembly. Evaluation of the pairwise registration is performed on the TUD-L [18] and ModelNet-40 [19] datasets. We follow [22] for prepossessing the TUD-L dataset. We split all datasets into train/val/test sets following existing literature for fair comparisons. These datasets define parts at distinct levels, ranging from random partitions (e.g., ModelNet-40 and BreakingBad) to human-labeled (e.g., semantically meaningful parts in PartNet and IKEA-Manual). The statistics and information of all datasets are summarized in Tab. 1.
@@
Following [11], we select the largest-volume part as the anchor and fix it during inference. However, this effectively provides the model with anchor pose in the object’s CoM (center of mass) frame, an unrealistic assumption for real-world assembly applications. Therefore, we also train our model in an anchor-free setting (see Appendix B: Anchor-free Models), and argue that anchor-free evaluation should be the standard protocol for shape assembly tasks.
@@
We evaluated our method against state-of-the-art methods for pairwise registration and shape assembly. For pairwise registration, we compare against DCPNet [31], RPMNet [30], GeoTransformer [32], and Diff-RPMNet [22]. For shape assembly, we compare against MSN [48], SE(3)-Assembly [49], Jigsaw [60], PuzzleFussion++ [25], and GARF [11]. We report our performances under two training configurations: dataset-specific training where models are trained independently for each dataset (denoted Ours (Single)), and joint training where a single model is trained across all datasets (denoted Ours (Joint)).
@@ 4.2 Evaluation
We report pose accuracy for shape assembly and pairwise registration in Tab. 2  [^1] We found that the BreakingBad benchmark [12, 11, 25] originally computed rotation error (RE) using the RMSE of Euler angles, which is not a proper metric on $\mathrm{SO(3)}$. To ensure consistency, we re-evaluate all baselines using the geodesic distance between rotation matrices via the Rodrigues formula [61, 31, 29, 32]. For Ours (Single) on TwoByTwo, we used encoder pretrained in the Ours (Joint) setting but trained the flow model only on TwoByTwo, due to its limited size. and Tab. 3, respectively. Our model outperforms all existing approaches by a substantial margin. For multi-part assembly, the closest competitor is GARF [11], which formulates per‐part pose estimation as 6-DoF pose regression; see Figs. 4 and 5. We attribute our superior results to two key advantages of Rectified Point Flow: (i) in contrast to our closest competitor GARF [11] which performs 6-DoF pose regression, our dense shape‐and‐pose parametrization helps the model learn better global shape prior and fine-grained geometric details more effectively; and (ii) our generative formulation natively handles part symmetries and interchangeability. For pairwise registration, GARF–despite being retrained on target datasets–fails to generalize beyond the original task. In contrast, our method achieves a new state‐of‐the‐art performance on registration benchmarks, outperforming methods designed solely for registration (e.g., GeoTransformer [32] and Diff-RPMNet [22]) and demonstrating strong generalization across different datasets (Fig 5). We also achieve the strongest shape assembly performance on IKEA-Manual [17]; for more details on evaluation and visualizations, see supplementary.
@@
Joint Training.
By recasting pairwise registration as a two-part assembly task, our unified formulation enables joint training of the flow model on all six datasets—including very small sets like TwoByTwo (308 samples) and IKEA-Manual (84 samples)—and the additional pretraining data from Objaverse. Ours (Joint) consistently matches or outperforms the dataset-specific (Ours (Single)) models. For example, on TwoByTwo the rotation error drops from 18.7 to 13.2 ($\approx 30\%$), and on BreakingBad from 9.6 to 7.4 ($\approx 23\%$), while on ModelNet-40, the rotation error is reduced from 1.37° to 0.93°. These results demonstrate that joint training enables the model to learn shared geometric priors from datasets with diverse part segmentation, symmetries, and common pose distributions, which substantially boosts performance, particularly on datasets with limited training samples.
@@
We demonstrate our model’s ability to handle symmetry (Sec. E) on IKEA-Manual [17], a dataset with symmetric assembly configurations. As shown in Fig. 7, while being only trained on a single configuration (left), Rectified Point Flow samples various reasonable assembly configurations (right), conditioned on the same input unposed point clouds. Note how our model permutes the 12 repetitive vertical columns and swaps the 2 middle baskets, yet always retains the non-interchangeable top and footed bottom baskets in their unique positions.
@@
Dataset Encoder Pre-training Task RE ↓ TE ↓ Part Acc ↑
@@
BreakingBad-Everyday [12] MLP No Pre-training 41.7 12.3 68.3
PTv3 [54] No Pre-training 18.5 4.9 79.5
PTv3 [54] Instance Segmentation 16.7 4.4 80.9
PTv3 [54] Overlap Detection (ours) 9.6 1.8 93.5
PartNet-Assembly [15] Point-BERT [62] Point Cloud Completion 27.4 23.3 45.2
PTv3 [54] Overlap Detection (ours) 24.8 15.4 50.2
Ablation on Overlap-aware Pretraining.
The first block of Tab. 4 compares four pretraining strategies for our flow‐based assembly model on BreakingBad–Everyday [12]. The first two encoders (MLP and PTv3 without pre-training) are trained jointly with the flow model. The last two encoders are PTv3 pretrained on instance segmentation and our overlap-aware prediction tasks, respectively. Their pretrained weights are frozen during flow model training. We find that PTv3 is a more powerful feature encoder compared to the MLP, and pretraining on instance segmentation can already extract useful features for pose estimation, while our proposed overlap‐aware pretraining leads to the best accuracy. We hypothesize that, although the segmentation backbone provides strong semantic features, only our overlap prediction task explicitly encourages the encoder to learn fine‐grained part interactions and pre‐assembly cues, critical for precise assembly and registration.
To further compare against autoencoder-based pretraining, we substitute our encoder with Point-BERT [62]’s encoder, which is pre-trained on ShapeNet [63], a superset of PartNet. We then train the flow model on PartNet under the same protocol. As reported in Tab. 4, Point-BERT encoder reduces the Part Accuracy from 50.2% to 45.2%, and worsens the accuracy in both TE and RE. We attribute this performance drop to: (i) Point-BERT is pre-trained for masked point cloud completion, which does not explicitly encourage the encoder to capture inter-part relations (e.g., contacts) that our overlap-aware pretraining targets; and (ii) PointBERT’s default 64-group tokenization aggregates points into relatively coarse groups, losing fine-grained geometric details for accurate pose estimation.
Shape Prior Learning.
Table 5: Part Accuracy [%] on Testing Part Schemes. Our model shows much less drop on OOD partitions.
Partition Scheme GARF Ours
Horizontal (ID) 99.5 100.0
Axial (OOD) 89.5 97.0
Random (OOD) 87.5 97.8
All 92.1 98.3
To probe whether our model learns the shape priors of assembled objects better than pose-vector methods, we construct a cylindrical toy dataset. We train using a single part partition scheme and then evaluate on the same cylinder shapes but with different partition schemes.
Specifically, we generated 6,000 training cylinders with heights and diameters uniformly sampled from $[0.2,1.0]$ m, each cut into two parts by a horizontal plane at a random height. For testing, 600 new cylinders were cut under three schemes: (i) Horizontal (in-distribution, ID), (ii) Axial: through the central axis at a random orientation (out-of-distribution, OOD), and (iii) Random: through random 3D plane (OOD). As shown in Tab. 5, our method achieved part accuracies of 100.0%, 97.0%, and 97.8% on three partition schemes, respectively. While GARF has comparable performance on the ID scheme, it degrades on two OOD schemes, verifying that our model learns a transferable shape prior over the whole assembled object rather than overfitting to the part partition-specific patterns.
Table 6: Zero-shot Evaluation on the Unseen Fractura Testset. We report Part Accuracy (%) for GARF and ours and include the supervised performance for reference.
Setting Method Leg Hip Rib Vertebra Pig Bones All
Supervised GARF 89.7 80.8 74.8 60.8 79.0 77.3
Zero-shot GARF 70.5 72.8 62.9 37.7 53.4 57.7
Ours 79.9 63.4 76.2 42.0 63.2 64.4
Generalization to Unseen Dataset.
We evaluated the out-of-domain generalization of our model by performing zero-shot tests on unseen bone fracture on the test split of the Fractura dataset [11], covering human bones (Leg, Hip, Vertebra, Rib) and pig bones. We report Part Accuracy for our method and GARF [11] in Tab. 6. Our model achieves strong zero-shot performance, surpassing GARF in most categories, especially Leg and Rib, and higher overall accuracy. While there remains a gap to fully supervised training, we expect that pretraining and/or fine-tuning on medical datasets will further improve our model by instilling bone-specific shape priors and fracture geometry cues.